# From Dataset to Real-world: General 3D Object Detection via Generalized Cross-domain Few-shot Learning

### v1

@@ 5.3 Ablation Studies
Table 5: Performance in mAP(%) with box searching by $L_{OD}$ only or $L_{box}$ (w/ $L_{OD}$, $L_{FVD}$, and $L_{BVC}$).
Box Search CS SS
$L_{OD}$ 4.65 12.56
$L_{box}$ 6.74 13.58
Component Ablation. Table 3 indicates that the proposed contrastive-enhanced prototype learning significantly enhances performance across overall classes with a notable boost for common classes, as (a)$\rightarrow$(b), illustrating its swift adaptation to limited samples in target domain. When applying our image-guided multi-modal fusion, the detection model shows marked improvement, especially on novel classes, as (a)$\rightarrow$(c), showing its boost on object recall. By combining both, our GCFS method achieves the highest performance, showing their complementary to each other.
We also conduct ablations on our proposed prototype learning and box searching components. Table 5 shows that the contrastive loss boosts model performance, indicating its ability to swiftly adapt prototypes to few-shot data in target domain. In Table 5, integrating $L_{OD}$ with $L_{FVD}$ and $L_{BVC}$ yields improvements on both CS and SS objects, showing $L_{FVD}$ and $L_{BVC}$ enhancing recall rates for objects with diverse structural complexities, thereby further optimizing model effectiveness.
Ablation on different shots. From Table 6, we observe, as K increases, our method demonstrates strong scalability to multi-shot settings, effectively improving both common and novel class detection. At K=40, our overall performance (26.05%) closely approaches full-shot (28.21%), bridging the gap between few-shot and full supervision. Notably, our novel-class performance surpasses full-shot training, due to data imbalance in full-shot learning and our image-guided approach enhancing novel object discovery. These results confirm that our method scales well with more shots, achieving strong generalization even with limited data.
@@
Table 6: Performances in mAP(%) with different $K$. (Full-shot denotes the training on the complete KITTI train set under pre-training settings.)
K Common Novel Overall
1 7.27 0.57 3.44
3 12.27 7.76 9.70
5 15.99 11.72 13.55
10 23.56 12.21 17.08
20 27.59 17.49 21.82
40 32.05 21.55 26.05
Full-shot 41.34 18.35 28.21
Table 7: Comparison in mAP(%) with OV and DA methods in the unsupervised few-shot setting. (Note that Unsup3D [24] and OpenSight [51] require sequential point clouds as inputs, which doesn’t apply to the discrete few-shot target data in GCFS.)
Methods Venue Common Novel Overall
DA SN [40] CVPR’20 12.09 - -
ST3D++ [48, 47] PAMI’22 21.00 - -
OVD FnP [6] ECCV’24 10.59 2.66 6.06
Ours-OVD - 22.25 8.26 14.26
@@
Unsupervised few-shot ablation with OV and DA methods. We establish an unsupervised few-shot setting with no annotations for all classes. Our approach is benchmarked against SOTA 3D-OVD solution, FnP [6], and well-established 3D-DA methods, SN [40] and ST3D++ [48, 47]. To create an OVD version of our method, we incorporate a physical-aware box searcher to generate high-quality pseudo-labels for target-specific training. Table 7 shows that, compared to OVD and DA methods, our OVD method achieves the highest mAPs, showing strong domain gap bridging capability and high learning efficiency from unlabeled samples. Please refer to the supplementary for implementation and result details.
@@ 6 Conclusion
This paper proposed the generalized cross-domain few-shot task in 3D object detection, aiming to adapt a source-pretrained model for strong performance on common and novel classes in the target domain with few-shot samples. We introduced the first GCFS solution, featuring a multi-modal fusion module for feature fusion and box proposal searching and contrastive-enhanced prototype learning for quick feature adaptation. In the established GCFS benchmarking settings, extensive experiments validated the effectiveness of our solution.
Supplementary Material
In this supplementary material, we provide additional details on our proposed image-guided multi-modal fusion module and the optimization-based meta-learning scheme in Section 7. In addition, we provide further details on the benchmark settings and method implementations, along with a comprehensive presentation of the experimental results in Section 8.
7 Details on Methodology
7.1 Image-guided Multi-modal Fusion
Front-viewed boundary calculation. Regarding the FVD loss, we utilize the LiDAR’s view angle to obtain the front-reviewed boundary FB for each class. Specifically, for given box $[x,y,z,\theta]$, we first get the yaw angle by $\alpha=\arctan(\frac{x}{y})$. Then, we obtain the view angle $\phi=\alpha-\theta$, which indicates the LiDAR scanning direction w.r.t. the search box. Regarding $\phi$, we define FB via the prior box length $l$ and width $w$:
@@ 7.1 Image-guided Multi-modal Fusion
7.2 Optimization-based Meta-learning
In meta-training, we utilize the sufficient data of common classes $C_{com}$ and other classes $C_{bas}^{s}$ to simulate the target few-shot fine-tuning on $C_{com}$ and $C_{nov}^{t}$. Specifically, we first randomly sample $N_{nov}$ classes from $C_{bas}^{s}$, where $N_{nov}$ is the class number of $C_{nov}^{t}$. Then, as shown in Figure 5, in the inner loop, we set up a $K$-shot cross-domain detection task, covering common classes and sampled $N_{nov}$ classes. For the meta-task, we utilize data augmentations on support data to simulate domain gaps regarding source-trained prototypes transferred to target data. Liked in MAML [8], we design the outer loop, where we run one inner loop to get the one-step updated parameter, run another inner loop to get the two-step gradient, and use the two-step gradient to update the original model parameters. As in [8], this inner-outer-loop meta-learning will find a set of parameters and prototypes $\textbf{F}_{proto}$ with the quick adaptation to a new few-shot learning task (covering common classes and novel classes) in a new domain via limited target data.
@@ 7.2 Optimization-based Meta-learning
8 Details on Experiments
8.1 Experiment Settings
@@
Implementation details. For the consistency of input point clouds across datasets, we unify the LiDAR coordinate system of all datasets by setting the origin on the ground. We adopt the point cloud range of [$-75.2m,-75.2m,-2m,75.2m,75.2m,4m$] and the voxel size of [$0.1m,0.1m,0.15m$]. For ground-truth sampling augmentation, we utilize its image-involved version implemented for KITTI in [35] and extend it to A2D2. Regarding class-specific attention module, the head number of is 4 and dropout rate is 0.1. During meta-training, we apply point density-resampling [17] on support data to enlarge domain shifts between query and support data. All experiments are conducted on 2× GeForce RTX-3090. Epoch numbers of the meta-training are limited 5 for NuScenes and Waymo, and 15 for KITTI. Batch sizes are 2 during pre-training and meta-training and 1 in few-shot fine-tuning and testing. $\lambda$, and $\lambda_{1}$, $\lambda_{2}$ are set to 1.0, 0.2, and 0.2.
@@
Compared methods. Indoor FSL methods (i.e., Proto-Vote [53], PVAE-Vote [37], and CP-Vote [18]), are mainly designed for the detection on novel classes, we extend their processing to common classes, to fit the GCFS tasks. Since Proto-Vote [53] is implemented for the indoor RGBD-based data with VoteNet [28] as the base detection model, we extend it to the outdoor LiDAR-abased data with the VoxelRCNN [5] as the base detection model, following our experiment setting. Considering no public codebase for PVAE-Vote and CP-Vote, we follow the paper methodologies and implementation details in papers and extend them to the GCFS tasks. For PVAE-Vote, given that the instability of VAE training is particularly pronounced in outdoor sparse and various point clouds, we incorporate the skip-connection architecture similar to ResNet, which enables VAE branches to learn residuals, thereby enhancing the stability of few-shot training. Regarding outdoor GFSL method GFS-Det with no public codebase, we follow the paper methodology and implementation details in [19] to extend it to the GCFS tasks.
@@ 8.2 Experimental Results
Tables 8, 9, 10, 11, 12, and 13, show more detailed performance among all classes w.r.t. Table 2. As shown in them, across all GCFS tasks, compared to existing methods, our method achieves more accurate object detection performance for overall common and novel classes, especially for “Car”, “Truck”, “Van”, “Cyclist”, “Tram”, “Utility_Vehicle”, and “Bus”. It demonstrates our method’s strong knowledge transferability from the common object in the source domain while effectively generalizing to novel classes with limited few-shot samples. Regarding the evaluation on high-density (64-beam) KITTI or low-density (16-beam) A2D2, the superior performance of our method underscores its strong adaptability to both moderate and extreme domain shifts. Especially in KITTI $\rightarrow 5$shot-A2D2 GCFS task, the performance of our method overpass the second-best performance significantly (i.e., overall mAP: VoxelRCNN $2.74\%\rightarrow 6.5\%$ and PVRCNN++ $3.17\%\rightarrow 7.76\%$), ensuring robust few-shot detection even in challenging low-density scenarios. The results of indoor 3D FSL methods (i.e., Proto-Vote, PVAE-Vote, CP-Vote) reflect the challenges in extending them to outdoor scenarios that is characterized by sparse point clouds at greater distances, dynamic objects, and varying lighting and weather conditions. Especially for common classes shared between source data and target data, those methods struggle with the demands of outdoor environments, resulting in reduced accuracy and robustness in the outdoor detection contexts. GFS-Det performs well mostly in common classes, especially on “Pedestrian” objects, indicating that its dedicated category-specific branches reduce interference of novel objects to common objects well learned in source pre-training. Yet, this separate-branch learning strategy forces the novel-object branch to learn geometric features from scratch, preventing it from leveraging geometric priors from common classes like cars or pedestrians. As a consequence, GFS-Det struggles with novel classes, hindering its ability to generalize effectively to newly learned objects and limiting its adaptability in GCFS tasks.
@@
Table 14 shows the performance of our proposed image-guided multi-modal fusion (denoted as Image-Fusion) method and our proposed contrastive-learning-enhanced prototype learning (denoted as CL-Proto) among all object classes. The experimental results show that our method demonstrates significant advantages in both novel and common classes, especially when combining Image-Fusion and CL-Proto, as in row (d), where it achieves the best performance. Specifically, the introduction of Image-Fusion significantly improves the performance on novel classes, raising the mAP from 5.48% to 11.17%. This improvement is particularly evident in classes like “Van”, “Cyclist”, and “Tram”, where data scarcity makes single-modal features insufficient. Leveraging image-guided multi-modal fusion enables the model to better capture features in novel classes, enhancing adaptability in few-shot scenarios. On the other hand, our proposed contrastive-learning-enhanced prototype learning mainly enhances the performance on common classes. When CL-Proto is added alone, the mAP for common classes increases from 12.77% to 14.80%, with a particularly notable improvement in the “Car” class, where mAP rises from 30.08% to 37.07%. Our proposed contrastive-learning-enhanced prototype learning improves the detection model with intra-class and inter-class differentiation, allowing the model to more accurately identify various features against source and target domain gaps. When Image-Fusion and CL-Proto are combined, as in row (d), the model achieves optimal performance in both novel and common classes, with an overall mAP reaching 13.55%. For novel classes, the mAP increases to 11.72%, and for common classes, it rises to 15.99%. This combination fully leverages the multi-modal feature representation strengths of our proposed image-guided multi-modal fusion method and our proposed contrastive-learning-enhanced prototype learning, enabling the model to perform better in the GCFS task.
@@
The results in Table 16 show that under the unsupervised few-shot setting, our extended OVD method demonstrates significant advantages across both common and novel classes. Although ST3D++ performs well in the car class, its performance is limited for other common classes, highlighting its lack of generalization in few-shot scenarios. Meanwhile, FnP’s initial advantage is largely due to its cautious box candidate search strategy, which works effectively in traditional OVD settings by leveraging a large amount of target data to accumulate good object samples. However, this approach is evidently inadequate for dealing with few-shot data, due to even fewer object samples for object feature learning. In contrast, our method achieves overall mAPs of 22.25% on common classes and 8.26% on novel classes, with a notable mAP boost on “Pedestrian”, “Van”, and “Cyclist”. This shows that our method, under the unsupervised few-shot setting, can effectively handle feature distribution differences in the target domain, achieving more accurate and balanced detection for both common and novel classes.

### v2

@@ 4.3 Ablation Studies
Table 4: Performance in mAP(%) of prototype learning with or without contrastive learning (CL).
Box Search CS $S$
$L_{\text{OD}}$ 4.65 12.56
$L_{\text{box}}$ 6.74 13.58
Table 5: Performance in mAP(%) with box searching by $L_{\text{OD}}$ only or $L_{\text{box}}$ (w/ $L_{\text{OD}}$, $L_{\text{FVD}}$, and $L_{\text{BVC}}$).
Methods
Target-
FT
Proto-
Vote
PVAE-
Vote
CP-Vote GFS Ours
Common 15.28 6.97 7.43 8.73 17.37 18.06
Novel 6.39 7.05 7.53 7.17 1.16 11.11
Overall 10.20 7.02 7.49 7.84 8.10 14.09
Table 6: Performance in mAP(%) of PV-RCNN for NuScenes $\rightarrow 5$shot-KITTI. Please refer to the supplementary for specifics across other GCFS tasks.
Shots K=1 K=3 K=5 K=10 K=20 K=40 Full-shot
Common 7.27 12.27 15.99 23.56 27.59 32.05 41.34
Novel 0.57 7.76 11.72 12.21 17.49 21.55 18.35
Overall 3.44 9.70 13.55 17.08 21.82 26.05 28.21
@@ Methods
Table 7: Performances in mAP(%) with different $K$. Full-shot denotes the training on the complete KITTI train set.
Method DA OVD
SN ST3D++ DenResamp FnP Ours-OVD
Common 12.09 21.00 14.89 10.59 22.25
Novel - - - 2.66 8.26
Overall - - - 6.06 14.26
Table 8: Comparison in mAP(%) for OVD and DA methods in the unsupervised few-shot setting.
[…]
We also conduct ablations on our proposed prototype learning and box searching components. Table 5 shows that the contrastive loss boosts model performance, indicating its ability to swiftly adapt prototypes to few-shot data in the target domain. In Table 5, integrating $L_{\text{OD}}$ with $L_{\text{FVD}}$ and $L_{\text{BVC}}$ yields improvements on both CS and SS objects, showing $L_{\text{FVD}}$ and $L_{\text{BVC}}$ enhancing recall rates for objects with diverse structural complexities, thereby further optimizing model effectiveness.
Ablation on Detection Backbone. We further evaluate our GCFS framework using the point-voxel-hybrid detector PV-RCNN. As shown in Table 6, our approach consistently outperforms others across all three metrics, demonstrating the generalizability of our solution.
Ablation on Numbers of Shots. Table 7 shows that our method scales well with increasing $K$. At $K=40$, the overall performance approaches the full-shot, narrowing the supervision gap. Despite a reasonable gap in common-class performance due to limited data, the novel-class performance surpasses the full-shot result, due to class imbalance in full-shot training and our image-guided design enhancing novel object discovery. These results confirm the scalability and generalization of our method under limited supervision.
@@
Unsupervised few-shot ablation with OVD and DA methods. We establish an unsupervised few-shot setting with no annotations for all classes. Our approach is benchmarked against the SOTA 3D-OVD solution and well-established 3D-DA methods in Table 8. To create an OVD version of our method, we incorporate a physical-aware box searcher to generate high-quality pseudo-labels for target-specific training. Compared to OVD and DA methods, our OVD method achieves the highest mAPs, showing strong domain gap bridging capability and high learning efficiency from unlabeled samples. Please refer to the supplementary for implementation and result details.
@@ 5 Conclusion
This paper tackled the generalized cross-domain few-shot task in 3D object detection and introduced the first GCFS solution. Beyond achieving state-of-the-art performance on four GCFS benchmarks, our work demonstrated a generalizable approach to few-shot 3D adaptation, grounded in the idea that semantic alignment across modalities and domains could be achieved by combining 2D open-set priors with 3D structural cues and few-shot supervision. We believed this framework opens new possibilities for 3D perception systems that must continually adapt to new environments and emerging object types, without relying on exhaustive data collection or domain-specific engineering.
Acknowledgements
This work was supported in part by JST CRONOS Grant (No. JPMJCS24K8), JSPS KAKENHI Grant (No.JP21H04877, No.JP23H03372, and No.JP24K02920), Canada CIFAR AI Chairs Program, the Natural Sciences and Engineering Research Council of Canada, and the Autoware Foundation.
Appendix A Details on Methodology
A.1 Image-guided Multi-modal Fusion
Box searching calculation. According to the center $[x,y,z]$ and heading angle $\theta$ of each object, the corresponding rotation transformation matrix T is defined as:
Regarding the FVD loss, we utilize the LiDAR’s view angle to obtain the front-reviewed boundary FB for each class. Specifically, for given box $x,y,z,\theta]$, we first get the yaw angle by $\alpha=\arctan(\frac{x}{y})$. Then, we obtain the view angle $\phi=\alpha-\theta$, which indicates the LiDAR scanning direction w.r.t. the search box. Regarding $\phi$, we define FB via the prior box length $l$ and width $w$:
@@ A.1 Image-guided Multi-modal Fusion
A.2 Optimization-based Meta-learning
In meta-training, we utilize the sufficient data of common classes $\mathbb{C}_{\text{com}}$ and other classes $\mathbb{C}_{\text{nov}}^{s}$ to simulate the target few-shot fine-tuning on $\mathbb{C}_{\text{com}}$ and $\mathbb{C}_{\text{nov}}^{t}$. Specifically, we first randomly sample $N_{\text{nov}}$ classes from $\mathbb{C}_{\text{nov}}^{s}$, where $N_{\text{nov}}$ is the class number of $\mathbb{C}_{\text{nov}}^{t}$. Then, as shown in Figure 5, in the inner loop, we set up a $K$-shot cross-domain detection task, covering common classes and sampled $N_{\text{nov}}$ classes. For the meta-task, we utilize data augmentations on support data to simulate domain gaps regarding source-trained prototypes transferred to target data. Like MAML (Finn et al. 2017), we design the outer loop, where we run one inner loop to get the one-step updated parameter, run another inner loop to get the two-step gradient, and use the two-step gradient to update the original model parameters. As in (Finn et al. 2017), this inner-outer-loop meta-learning will find a set of parameters and prototypes $\textbf{F}_{\text{proto}}$ with the quick adaptation to a new few-shot learning task (covering common classes and novel classes) in a new domain via limited target data.
@@ A.2 Optimization-based Meta-learning
Table 15: Performance comparison in mAP(%) on VoxelRCNN detector in the KITTI $\rightarrow 5$shot-Argo2 GCFS task. The bold values represent the best performance. Subscript values in parentheses are standard deviations. The mtc for Motorcycle, tcn is for Traffic_cone, lve for Large_vehicle, and cbl for Construction_barrel.
Methods Venues car ped trk common bcy bus mtc tcn lve cbl sign novel overall
Target-FT - 4.70(0.32) 1.66(0.12) 3.20(0.12) 3.18(0.19) 0.10(0.07) 2.96(0.51) 0.01(0.01) 0.03(0.01) 0.14(0.10) 0.12(0.09) 0.99(0.09) 0.62(0.12) 1.39(0.14)
Proto-Vote NIPS’22 4.16(1.45) 1.40(0.55) 4.42(0.68) 3.33(0.90) 1.98(0.87) 2.19(0.34) 0.49(0.04) 0.14(0.10) 0.38(0.39) 0.57(0.44) 0.53(0.36) 0.90(0.36) 1.63(0.52)
PVAE-Vote NIPS’24 4.64(1.92) 1.54(0.57) 3.11(0.61) 3.10(1.03) 1.21(0.14) 3.45(0.43) 0.23(0.04) 0.38(0.08) 0.63(0.53) 0.42(0.22) 0.14(0.39) 0.92(0.26) 1.58(0.49)
CP-Vote PRCV’24 3.92(1.52) 1.26(0.60) 3.00(0.48) 2.72(0.87) 1.15(0.58) 3.51(0.45) 0.16(0.05) 0.24(0.26) 0.27(0.18) 0.61(0.59) 0.55(0.41) 0.93(0.36) 1.47(0.51)
GFS arXiv’23 10.57(0.11) 4.82(0.09) 2.96(0.01) 6.11(0.07) 0.07(0.02) 0.02(0) 0.04(0.01) 0.08(0.03) 0.02(0.00) 0.00(0.00) 0.00(0.00) 0.03(0.01) 1.86(0.03)
Ours - 7.78(0.22) 4.38(0.12) 7.97(0.13) 6.71(0.16) 1.38(0.05) 8.25(0.98) 0.63(0.01) 0.58(0.03) 0.97(0.11) 0.70(0.04) 1.97(0.01) 2.07(0.17) 3.46(0.17)
Appendix B Details on Experiments
B.1 Experiment Settings
GCFS benchmark settings. Nuscenes (Caesar et al. 2020) contains labeled point cloud samples collected by a 32-beam LiDAR, mainly covering classes: Car, Pedestrian, Truck, Bicycle, Barrier, Construction_vehicle, Bus, Trailer, Motorcycle, Traffic_cone, etc. We use the train set ($\sim$28K samples) including 3 common classes for model pertaining and the train set including 3 common classes and the rest 7 classes for meta pertaining. Waymo (Sun et al. 2020) utilizes a 64-beam spinning-scanning LiDAR and 4 forward-scanning LiDARs to collect point cloud samples. The original Waymo contains 4 classes, i.e. {Vehicle, Pedestrian, Cyclist, Sign}, where “Vehicle” covers cars, trucks, buses, motorcycles, bicycles, etc. To align the class categories in cross-domain settings, we refine the “Vehicle” class labels into 5 different classes {Car, Truck, Bus, Motorcycle, Bicycle}. Specifically, with the help of finer-grained segmentation annotations (involving those 5 vehicle classes), we use the point-wise segmentation label to relabel vehicle objects into the object category with the highest number of points. For refined Waymo, we use the segmentation-involved $\sim$24K train samples, including 3 common classes for model pertaining and train samples including 3 common classes and 4 source-specific classes {Bus, Motorcycle, Bicycle, Sign} for meta pertaining. KITTI (Geiger et al. 2012) contains $\sim$7K labeled samples collected by a 64-beam LiDAR, covering classes: Car, Pedestrian, Truck, Van, Person_sitting, Cyclist, Tram. A2D2 (Geyer et al. 2020) contains $\sim$12K labeled samples collected by a 16-beam LiDAR, mainly covering classes: Car, Pedestrian, Truck, Bicycle, Utility_vehicle, Bus, etc. Argoverse 2 (Wilson et al. 2023) provides 1000 scenes and $\sim$134K labeled LiDAR frames collected by the top-mounted 32-beam spinning LiDAR, covering a wide variety of urban scenes and diverse object categories such as Vehicle, Pedestrian, Bus, Bicycle, Motorcycle, Construction_barrel, and Construction_vehicle.
For FS-KITTI, we randomly select samples from the train set to form the training data, including $K$-shot objects for each class, and use the complete val set for model evaluation. For FS-A2D2, we first randomly select 6 out of 12 sequences and randomly select samples to form the training data, including $K$-shot objects for each class. We take the other 6 sequences as test data for model evaluation. Since the samples of FS-A2D2 are temporarily sequential, we uniformly sample 50% of test data for computational efficiency during model evaluation. Also, we remove a small number of test samples with erroneous annotations (e.g., labeling “Car” as “Pedestrian”) and ambiguous class labels (e.g., labeling “Bus” as “Truck”) for a fair evaluation. Code implementation and more detailed information on the data set split are available in our codebase.
@@ B.1 Experiment Settings
Implementation details. For the consistency of input point clouds across datasets, we unify the LiDAR coordinate system of all datasets by setting the origin on the ground. We adopt the point cloud range of [75.2m,75.2m,2m,75.2m,75.2m,4m-75.2m,-75.2m,-2m,75.2m,75.2m,4m] and the voxel size of [$0.1m,0.1m,0.15m$]. For ground-truth sampling augmentation, we utilize its image-involved version implemented for KITTI in (Song et al. 2024) and extend it to A2D2. Regarding the class-specific attention module, the head number is 4, and the dropout rate is 0.1. During meta-training, we apply point density-resampling (Li et al. 2025) on support data to enlarge domain shifts between query and support data. All experiments are conducted on 2× GeForce RTX-3090 with a total memory of 48GB. Our code implementation is based on the codebase of OpenPCDet (Team 2020) and RoboFusion (Song et al. 2024). For full-shot target learning, we use the training setting the same as the pre-training setting (e.g., augmentation, learning rate, optimizer, hyper-parameters) as in Section 4, and the epoch numbers of the full-shot training epoch are 80 for KITTI and A2D2, and 6 for Argoverse 2. For meta-training, epoch numbers are limited to 5 for NuScenes and Waymo, and 15 for KITTI, to obtain swiftly-adaptive model weights. Batch sizes are 2 during pre-training and meta-training and 1 in few-shot fine-tuning and testing. $\lambda$, and $\lambda_{1}$, $\lambda_{2}$ are set to 1.0, 0.2, and 0.2. The temperature $\tau$ in the InfoNCE loss is set to 0.07.
@@
Compared methods. Indoor FSL methods (i.e., Proto-Vote (Zhao and Qi 2022), PVAE-Vote (Tang et al. 2024), and CP-Vote (Li et al. 2024)) are mainly designed for the detection of novel classes. We extend their processing to common classes to fit the GCFS tasks. Since Proto-Vote (Zhao and Qi 2022) is implemented for the indoor RGBD-based data with VoteNet (Qi et al. 2019) as the base detection model, we extend it to the outdoor LiDAR-based data with the VoxelRCNN (Deng et al. 2021) as the base detection model, following our experiment setting. Considering no public codebase for PVAE-Vote and CP-Vote, we follow the paper methodologies and implementation details in the papers and extend them to the GCFS tasks. For PVAE-Vote, given that the instability of VAE training is particularly pronounced in outdoor sparse and various point clouds, we incorporate the skip-connection architecture similar to ResNet, which enables VAE branches to learn residuals, thereby enhancing the stability of few-shot training. Regarding outdoor GFSL method GFS-Det with no public codebase, we follow the paper methodology and implementation details in (Liu et al. 2023) to extend it to the GCFS tasks. DenResamp  (Li et al. 2025) proposes a single-domain generalization method that utilizes density-resampling-based augmentation and test-time adaptation to bridge density-related domain gaps. We explore its domain-adaptive version developed in the paper (Li et al. 2025) as a 3D-DA method.
@@ B.2 Experimental Results
Tables 9 to 15 show more detailed performance among all classes. As shown in them, across all GCFS tasks, compared to existing methods, our method achieves more accurate object detection performance for overall common and novel classes, especially for “Car”, “Truck”, “Van”, “Cyclist”, “Tram”, “Utility_Vehicle”, and “Bus”. It demonstrates our method’s strong knowledge transferability from the common object in the source domain while effectively generalizing to novel classes with the few-shot samples. Regarding the evaluation on high-density (64-beam) KITTI or low-density (16-beam) A2D2 or (32-beam) Argoverse 2, the superior performance of our method underscores its strong adaptability to both moderate and extreme domain shifts. Especially in KITTI $\rightarrow 5$shot-A2D2 GCFS task (Tables 13 and 14), the performance of our method surpasses the second-best performance significantly (i.e., overall mAP: VoxelRCNN $2.74\%\rightarrow 6.5\%$ and PVRCNN++ $3.17\%\rightarrow 7.76\%$), ensuring robust few-shot detection even in challenging low-density scenarios. Also, as shown in Table 15, where an extreme semantic shift exists, our proposed method shows the highest overall performance, indicating its superiority on fast adaptation to novel semantics under minimal target supervision. The results of indoor 3D FSL methods (i.e., Proto-Vote, PVAE-Vote, CP-Vote) reflect the challenges in extending them to outdoor scenarios that are characterized by sparse point clouds at greater distances, dynamic objects, and varying lighting and weather conditions. Especially for common classes shared between source data and target data, those methods struggle with the demands of outdoor environments, resulting in reduced accuracy and robustness in the outdoor detection contexts. GFS-Det performs well mostly in common classes, especially on “Pedestrian” objects, indicating that its dedicated category-specific branches reduce the interference of novel objects to common objects well learned in source pre-training. Yet, this separate-branch learning strategy forces the novel-object branch to learn geometric features from scratch, preventing it from leveraging geometric priors from common classes like cars or pedestrians. As a consequence, GFS-Det struggles with novel classes, hindering its ability to generalize effectively to newly learned objects and limiting its adaptability in GCFS tasks.
@@
Table 16 shows the performance of our proposed image-guided multi-modal fusion (denoted as Image-Fusion) method and our proposed contrastive-learning-enhanced prototype learning (denoted as CL-Proto) among all object classes. The experimental results show that our method demonstrates significant advantages in both novel and common classes, especially when combining Image-Fusion and CL-Proto, as in row (d), where it achieves the best performance. Specifically, the introduction of Image-Fusion significantly improves the performance on novel classes, raising the mAP from 5.48% to 11.17%. This improvement is particularly evident in classes like “Van”, “Cyclist”, and “Tram”, where data scarcity makes single-modal features insufficient. Leveraging image-guided multi-modal fusion enables the model to better capture features in novel classes, enhancing adaptability in few-shot scenarios. On the other hand, our proposed contrastive-learning-enhanced prototype learning mainly enhances the performance on common classes. When CL-Proto is added alone, the mAP for common classes increases from 12.77% to 14.80%, with a particularly notable improvement in the “Car” class, where mAP rises from 30.08% to 37.07%. Our proposed contrastive-learning-enhanced prototype learning improves the detection model with intra-class and inter-class differentiation, allowing the model to more accurately identify various features against source and target domain gaps. When Image-Fusion and CL-Proto are combined, as in row (d), the model achieves optimal performance in both novel and common classes, with an overall mAP reaching 13.55%. For novel classes, the mAP increases to 11.72%, and for common classes, it rises to 15.99%. This combination fully leverages the multi-modal feature representation strengths of our proposed image-guided multi-modal fusion method and our proposed contrastive-learning-enhanced prototype learning, enabling the model to perform better in the GCFS task. Notably, through mixed-precision VLM acceleration, optimizations to the model’s pre- and post-processing, and other engineering improvements, our method achieves 10.11 FPS on the NVIDIA A100 GPU in the representative NuScenes→KITTI setting.
@@
The results in Table 18 show that under the unsupervised few-shot setting, our extended OVD method demonstrates significant advantages across both common and novel classes. Although ST3D++ performs well in the car class, its performance is limited for other common classes, highlighting its lack of generalization in few-shot scenarios. Meanwhile, FnP’s initial advantage is largely due to its cautious box candidate search strategy, which works effectively in traditional OVD settings by leveraging a large amount of target data to accumulate good object samples. However, this approach is inadequate for dealing with few-shot data, due to even fewer object samples for object feature learning. In contrast, our method achieves overall mAPs of 22.25% on common classes and 8.26% on novel classes, with a notable mAP boost on “Pedestrian”, “Van”, and “Cyclist”. This shows that our method, under the unsupervised few-shot setting, can effectively handle feature distribution differences in the target domain, achieving more accurate and balanced detection for both common and novel classes.
Table 19: BEV/FV AP (%) in the bird’s eye view and front view of the VoxelRCNN detection under NuScenes $\rightarrow 5$shot-KITTI (N→FS-K), Waymo $\rightarrow 5$shot-KITTI (W→FS-K), KITTI $\rightarrow 5$shot-A2D2 (K→FS-A)
Settings Common Novel Overall
N→FS-K Target-FT 24.16/36.16 6.00/7.32 13.79/19.68
Ours 29.39/40.57 12.7/16.37 19.86/26.74
W→FS-K Target-FT 31.89/47.12 13.54/15.56 21.4/29.08
Ours 33.89/48.84 18.53/22.23 25.11/33.63
K→FS-A Target-FT 10.1/45.92 0.86/4.05 5.48/24.99
Ours 16.03/61.33 5.73/11.03 10.88/36.18
Table 19 presents the 2D Average Precision (AP) results of VoxelRCNN under various cross-domain 5-shot settings. Compared to the Target-FT baseline, the proposed method consistently improves performance across all scenarios, especially for novel classes and in more challenging domain shifts such as KITTI to A2D2. Notably, the proposed method significantly boosts AP in both bird’s eye view (BEV) and front view (FV), demonstrating its strong generalization ability in few-shot settings. These improvements highlight the method’s effectiveness in enhancing detection for unseen categories and its robustness in handling domain discrepancies.